OpenStudy (anonymous):

What is the ratio 15 : 24 written in lowest terms? A. 3 : 5 B. 5 : 8 C. 3 : 4 D. 7 : 8

OpenStudy (igreen):

Divide both of the numbers by 3.

OpenStudy (anonymous):

\[15:24=\frac{15}{24}=\frac{5 \times 3}{8 \times 3}=\frac{?}{?}=?:?\]

OpenStudy (igreen):

15 / 3 = ? 24 / 3 = ?

OpenStudy (anonymous):

When you divide 15 by 3 you get 5 and when you divide 24 by 3 you get 8, so you get 5/8, which is B

OpenStudy (anonymous):

Also this is a good website to help you with math http://www.wolframalpha.com

OpenStudy (anonymous):

spry notifications didn't tell me.

OpenStudy (anonymous):

It is ok no problem

OpenStudy (anonymous):

Thanks for the website

OpenStudy (anonymous):

you are very welcome